OpenStudy (kimo_indie):

Solve the system of linear equations below. x - 3y = -3 x + 3y = 9 a) x = -12, y = 7 b) x = 3, y = 2 c) x = 6, y = 1 d) x = 6, y = 2

OpenStudy (rushwr):

Add both equation together. x - 3y + x + 3y = -3 +9 2x = 6 x = 3 Now substitute x = 3 in to any of the equations given I will substitute it to the 2nd equation x + 3y = 9 3 + 3y = 9 3y = 9-3 3y = 6 y = 2

OpenStudy (kimo_indie):

thanks so much!